Suppose that a company wishes to predict sales volume based on the amount of advertising expenditures. The sales manager thinks that sales volume and advertising expenditures are modeled according to the following linear equation. Both sales volume and advertising expenditures are in thousands of dollars.

Estimated Sales Volume=45.05+0.49(Advertising Expenditures)


If the company has a target sales volume of $275,000, how much should the sales manager allocate for advertising in the budget? Round your answer to the nearest dollar.
Answer #1

Estimated Sales Volume=45.05+0.49(Advertising Expenditures)

sales volume = 275000

275000 = =45.05+0.49(Advertising Expenditures)


Advertising Expenditures = 274954.95 / 0.49

= 561133